OpenStudy (lgg23):

Determine whether the sequence is divergent or convergent. If it is convergent, evaluate it's limit. If it is divergent, is it positive or negative infinity. A sub n= (13(3^n)+19)/(17(5^n))

OpenStudy (anonymous):

It appears to converge and the limit is 0. Think about it this way, for larger values of n, the constans won't matter, so we are left with:\[\lim_{n \rightarrow \infty} \frac{3^{n}}{5^n} = 0\]

OpenStudy (lgg23):

Oh okay. So I can apply some of the basics of limits to sequences.